The manager of a grocery store in the retirement community of Kapiri is interested in providing good service to the senior citizens who shop in his store. Presently, the store has a separate check-out counter for senior citizens. On average, 30 senior citizens per hour arrive at the counter according to a Poisson distribution and are served at an average rate of 35 customers per hour with exponential service times. Find the following;

(a) Utilization of the checkout clerk

(b) Number of customers in a system

(c) Number of customers in line

(d) Time spent in the system

(e) Waiting time in line

The manager of Kapiri grocery wants to answer the following question:

(1) What service rate would be required to have customers average only eight minutes in


1
Expert's answer
2021-05-03T08:32:24-0400

We have given,

L = Average arrival time = "\lambda" = 30 per hour

M = Average service rate = "\mu" = 35 per hour

a.) Utilization of the checkout clerk, "U = \dfrac{L}{M}" = "\dfrac{30}{35}" = "85.7 %" %

b.) Number of customers in a system,

Firstly we have to find out the Average number of customers or units waiting in line for service

"L_q = \dfrac{\lambda^2}{\mu(\mu-\lambda)}"


"= \dfrac{30\times 30}{35(35-30)}"


"L_q= 5.14" ,


Now, Average number of customers or units in the system = "L_q + \dfrac{\lambda}{\mu}"


= "5.14+\dfrac{30}{35}"

"= 5.997"

c.) Number of customers in a line "= L_q = 5.14"

d.) Time spent in the system

But before that we have to find out,

The Average time a customer or unit spends waiting in line for service "W_q = \dfrac{L_q}{\lambda}"

Hence, "W_q = \dfrac{5.14}{30} = 0.17"

Now, time spent in the system "W = W_q + \dfrac{1}{\mu}"


= "0.17 + \dfrac{1}{35}"


"= 0.19"

e.) Waiting time in line "= W_q = 0.17"


1.) Service rate would be required to have customers average only eight minutes in = "\dfrac{60}{8} \times 35 = 262.5"